sqrt(x+3)+1=sqrt(3x-2) Solve for x
Answer by mananth(16946)

sqrt(x+3)+1=sqrt(3x-2) Solve for x
...
square both sides
..
(sqrt(x+3)+1)^2=(sqrt(3x-2))^2
(x+3)+2sqrt(x+3)+1 = 3x-2
x+3 -3x+2+1= -2sqrt(x+3)
-2x+6=-2sqrt(x+3)
-2(x-3)=-2(sqrt(x+3)
divide by -2
(x-3)^2= sqrt(x+3)
square
x^2-6x+9 = x+3
x^2-6x-x+9-3=0
x^2-7x+6=0
x^2-6x-x+6=0
x(x-6)-1(x-6)=0
(x-6)(x-1)=0
x 6 Or 1
...
m.ananth@hotmail.ca